OpenStudy (anonymous):

Explain the role of stress and modern lifestyles in body and movement problems?

OpenStudy (anonymous):

\[stress = \frac{ 1 }{ movement} \] \[body = Cmovement\] where C is an arbitrary amount of will \[Lifestyle _{modern}= stress(neighbours Money - yourMoney)\] so \[stress + Lifestyle_{modern} = \frac{1}{Cmovement} + stress(neighboursMoney - yourMoney)\]

OpenStudy (anonymous):

\[stress + Life_{modern} = \frac{1}{cMove} +stress(neighbour$ - your$)\]

OpenStudy (anonymous):

no? :(

OpenStudy (anonymous):

Lol I don't get what this means ..like in sentence form?

OpenStudy (anonymous):

ohhh, im rubbish at sentence's

OpenStudy (fellowroot):

Stress on the body happens all the time. Take for example a person running, when their feet press down on the ground this causes stress in their knees and in all of their joints.